# Gravitational waves from burdened primordial black holes dark matter

### v1

@@ III.1 Scalar-induced gravitational wave from inflation
Using Eqs. 10 and 11, we compute the GW signal as a function of frequency for different PBH’s masses. We choose $f_{\rm PBH}=1$ so that PBHs can be the entire DM. In the high-frequency regime, there are some upcoming experiments that could be relevant for us such as LISA LISA:2017pwj , BBO Yagi:2011wg , DECIGO Kawamura:2011zz , CE LIGOScientific:2016wof , ET Punturo:2010zz . The frequency range and sensitivity of DECIGO/CE are similar to that of BBO/ET respectively, so we do not show them here to avoid crowded figures. We show our results together with the sensitivity curves of some notable GW experiments in Fig. 1.
@@
III.2 Gravitational wave from the first-order phase transition
In a FOPT, the energy gained in the transition from the false to the true vacuum is transferred into the kinetic energy of the bubble walls. When the bubbles expand and eventually collide, they release a huge amount of energy and some of them is in the form of GW. At the same time, the bubbles also interact with the cosmic fluid and, as they collide, they also produce turbulence and acoustic waves. In summary, GW from a FOPT can come from three different sources: bubble collision, sound wave, and turbulence Maggiore:2018sht . In many cases, the contribution from sound wave is most dominant. The total fractional energy density of GW today is
@@ III.2 Gravitational wave from the first-order phase transition
In the following, we mostly follow Caprini:2015zlo to have the form of GW spectrum from three different sources. By using Eqs. 18 and 19, we can conveniently express these contributions in terms of the PBH’s mass and population. This will provide a more direct correlation between the GW signal and the corresponding PBH’s features. It will also help us choose the best benchmark parameters for experimental prospects as well as make clearer comparison with the GW signal from inflation.
@@
$\Omega_{\rm col}$ $\approx 3.72\times 10^{-5}\left(\frac{100}{g_{*}}\right)^{1/3}\left(\frac{\beta}{H_{*}}\right)^{-2}\left(\frac{\kappa_{\phi}\alpha_{*}}{1+\alpha_{*}}\right)^{2}\frac{0.11v_{w}^{3}}{0.42+v_{w}^{3}}S_{\rm col}(f)$ (29)
$\approx 1.35\times 10^{-8}\alpha_{*}^{-1/2}\eta_{\chi}^{-2}\frac{v_{w}}{0.42+v_{w}^{2}}\left(\frac{\kappa_{\phi}\alpha_{*}}{1+\alpha_{*}}\right)^{2}\left(\frac{g_{*}}{100}\right)^{-1/2}\beta_{\rm PBH,i}^{\prime}S_{\rm col}(f),$ (30)
@@
The fraction of latent heat converted into the kinetic energy of the bubble walls is
@@
$\approx 3.25\times 10^{-7}\alpha_{*}^{-1/4}\eta_{\chi}^{-1}\left(\frac{\kappa_{\rm sw}\alpha_{*}}{1+\alpha_{*}}\right)^{2}\left(\frac{g_{*}}{100}\right)^{-5/12}(\beta_{\rm PBH,i}^{\prime})^{2/3}S_{\rm sw}(f)\Upsilon,$ (36)
@@
The fraction of latent heat converted into the bulk motion of plasma is
@@
$\approx 4.1\times 10^{-5}\alpha_{*}^{-1/4}\eta_{\chi}^{-1}\left(\frac{\kappa_{\rm turb}\alpha_{*}}{1+\alpha_{*}}\right)^{3/2}\left(\frac{g_{*}}{100}\right)^{-5/12}(\beta_{\rm PBH,i}^{\prime})^{2/3}S_{\rm turb}(f),$ (43)
@@
the fraction of latent heat converted into plasma turbulence is $\kappa_{\rm turb}\approx 0.1\kappa_{\rm sw}$ Borah:2024lml , and the frequency at the phase transition time is
@@
From the above expressions, we see that the GW signal depends only weakly on $v_{w}$. Therefore, let us fix this parameter with its typical value: $v_{w}\sim 0.6$ Moore:1995ua . The GW signal is more sensitive to $\alpha_{*}$ which typically can be of order $O(0.01-1)$. Although $\alpha_{*}\sim 1$ would give the strongest signal, let us be a bit conservative and pick $\alpha_{*}\sim 0.1$ as it is typically realized in many models Caprini:2015zlo . The peak of GW spectrum also depends on the asymmetry parameter $\eta_{\chi}$ and the PBH’s initial population $\beta_{\rm PBH,i}^{\prime}$. The latter can be fixed for a chosen mass of PBH and when we demand that PBHs must be entire DM with $f_{\rm PBH}=1$ (see Eq. 22). Therefore, given the fact that $\Omega_{\rm GW}\propto\eta_{\chi}^{-1}$, the smallest $\eta_{\chi}$ would give the strongest GW signal that we hope for. The actual value of $\eta_{\chi}$ depends on the underlying model and is generally unconstrained Hong:2020est  [^3] By definition, $\eta_{\chi}\lesssim n_{\chi}^{\rm eq}(T)/s(T)\sim 10^{-3}$ Kawana:2021tde , but there is no lower bound. The baryon asymmetry is $\sim 10^{-10}$ but $\chi$ could be in the dark sector.. We find that $\eta_{\chi}=\{10^{-16},10^{-15},10^{-14}\}$ would give $T_{*}\sim\{10^{7}\ \rm GeV,10^{6}\ \rm GeV,10^{5}\ \rm GeV\}$ and $\beta/H_{*}\sim 1$ for $M_{\rm PBH}=\{10^{3}\ \rm g,10^{6}\ \rm g,10^{9}\ \rm g\}$ respectively. Reducing further the value of $\eta_{\chi}$ would result in too small $\beta/H_{*}$ that is not viable in most scenarios of FOPT. Therefore, the mentioned set of $\eta_{\chi}$ is a good benchmark point as it gives the strongest possible GW signal. The total GW spectrum from the FOPT is shown in Fig. 2 for different PBH’s masses and $f_{\rm PBH}=1$. For an example of individual contributions to the GW spectrum, see Appendix B.
@@
  • The peak frequency is inversely proportional to the mass, which is similar to the inflation case. However, unlike inflation, it also depends on the PBH’s final population (see Eqs. 39 and 22). We observe that, although the magnitude of the GW peak depends on parameter choice of the underlying FOPT, the peak frequency is a robust prediction when we look for a specific mass and population of PBH.
@@
In the left panel of Fig. 3, we show $\Omega_{\rm GW,peak}$ as a function of $\eta_{\chi}$ for different PBH masses. Other fixed benchmark parameters are the same as before and are shown on the top of the panel. The light portion of the curves cannot be detected by current proposed experiments [^4] These just include some major proposals such as LISA, BBO, ET. Of course, if it happens that we can have more experiments with greater sensitivity, then more regions of the parameter space can be probed.. For each value of $\eta_{\chi}$ from the left panel, we can infer the corresponding values of $\beta/H_{*}$ and $T_{*}$, as shown in the right panel of Fig. 3. Therefore, a detection of such GW could be an indication of burdened PBHs as DM scenario and could also offer insights into the underlying details of FOPT.
@@ IV Conclusion
Appendix A: Redshift of GW’s spectrum and frequency in generic cosmological backgrounds

### v2

@@ III.1 Scalar-induced GW from inflation
Using Eqs. 11 and 12, we compute the GW signal as a function of frequency for different PBH’s masses. We choose $f_{\rm PBH}=1$ so that PBHs can be the entire DM. In the high-frequency regime, there are some upcoming experiments that could be relevant for us such as LISA LISA:2017pwj , BBO Yagi:2011wg , DECIGO Kawamura:2011zz , CE LIGOScientific:2016wof , ET Punturo:2010zz . The frequency range and sensitivity of DECIGO/CE are similar to that of BBO/ET respectively, so we do not show them here to avoid crowded figures. We show our results together with the sensitivity curves of some notable GW experiments [^6] The GW sensitivity curves of LISA and BBO were taken from Moore:2014lga , and that of ET was taken from ET:2019dnz . in Fig. 1.
@@
III.2 GW from FOPT
In a FOPT, the energy gained in the transition from the false to the true vacuum is transferred into the kinetic energy of the bubble walls. When the bubbles expand and eventually collide, they release a huge amount of energy and some of them is in the form of GW. At the same time, the bubbles also interact with the cosmic fluid and, as they collide, they also produce turbulence and acoustic waves. In summary, GW from a FOPT can come from three different sources: bubble collision, sound wave, and turbulence Maggiore:2018sht . Typically, the contribution from sound wave is most dominant Hindmarsh:2015qta . The total fractional energy density of GW today is
@@ III.2 GW from FOPT
In the following, we mostly follow Caprini:2015zlo to have the form of GW spectrum from three different sources. By using Eqs. 19 and 20, we can conveniently express these contributions in terms of the PBH’s mass and population. This will provide a more direct correlation between the GW signal and the corresponding PBH’s features. It will also help us choose the best benchmark parameters for experimental prospects as well as make clearer comparison with the GW signal from inflation.
@@
$\Omega_{\rm col}$ $\approx 3.72\times 10^{-5}\left(\frac{100}{g_{*}}\right)^{1/3}\left(\frac{\beta}{H_{*}}\right)^{-2}\left(\frac{\kappa_{\rm col}\alpha_{*}}{1+\alpha_{*}}\right)^{2}\frac{0.11v_{w}^{3}}{0.42+v_{w}^{2}}S_{\rm col}(f)$ (30)
$\approx 1.47\times 10^{-8}\alpha_{*}^{-1/2}\eta_{\chi}^{-2}\frac{v_{w}}{0.42+v_{w}^{2}}\left(\frac{\kappa_{\rm col}\alpha_{*}}{1+\alpha_{*}}\right)^{2}\left(\frac{g_{*}}{100}\right)^{-1/2}(\beta_{\rm PBH,i}^{\prime})^{4/3}S_{\rm col}(f),$ (31)
@@
The fraction of latent heat converted into the kinetic energy of the bubble walls is Kamionkowski:1993fg
@@
$\approx 3.54\times 10^{-7}\alpha_{*}^{-1/4}\eta_{\chi}^{-1}\left(\frac{\kappa_{\rm sw}\alpha_{*}}{1+\alpha_{*}}\right)^{2}\left(\frac{g_{*}}{100}\right)^{-5/12}(\beta_{\rm PBH,i}^{\prime})^{2/3}S_{\rm sw}(f)\Upsilon,$ (37)
@@
The fraction of latent heat converted into the bulk motion of plasma is Espinosa:2010hh
@@
$\approx 4.48\times 10^{-5}\alpha_{*}^{-1/4}\eta_{\chi}^{-1}\left(\frac{\kappa_{\rm turb}\alpha_{*}}{1+\alpha_{*}}\right)^{3/2}\left(\frac{g_{*}}{100}\right)^{-5/12}(\beta_{\rm PBH,i}^{\prime})^{2/3}S_{\rm turb}(f),$ (44)
@@
the fraction of latent heat converted into plasma turbulence is $\kappa_{\rm turb}\approx 0.1\kappa_{\rm sw}$ Borah:2024lml ; Hindmarsh:2015qta , and the frequency at the phase transition time is
@@
From the above expressions, we see that the GW signal depends only weakly on $v_{w}$. Therefore, let us fix this parameter with its typical value: $v_{w}\sim 0.6$ Moore:1995ua . The GW signal is more sensitive to $\alpha_{*}$ which typically can be of order $O(0.01-1)$. Although $\alpha_{*}\sim 1$ would give the strongest signal, let us be a bit conservative and pick $\alpha_{*}\sim 0.1$ as it is typically realized in many models Caprini:2015zlo . The peak of GW spectrum also depends on the asymmetry parameter $\eta_{\chi}$ and the PBH’s initial population $\beta_{\rm PBH,i}^{\prime}$. The latter can be fixed for a chosen mass of PBH and when we demand that PBHs must be entire DM with $f_{\rm PBH}=1$ (see Eq. 23). Therefore, given the fact that $\Omega_{\rm GW}\propto\eta_{\chi}^{-1}$, the smallest $\eta_{\chi}$ would give the strongest GW signal that we hope for. The actual value of $\eta_{\chi}$ depends on the underlying model and is generally unconstrained Hong:2020est  [^7] By definition, $\eta_{\chi}\lesssim n_{\chi}^{\rm eq}(T)/s(T)\sim 10^{-3}$ Kawana:2021tde , but there is no lower bound. The baryon asymmetry is $\sim 10^{-10}$ but $\chi$ is in the dark sector.. We find that $\eta_{\chi}=\{0.5\times 10^{-16},0.5\times 10^{-15},0.5\times 10^{-14}\}$ would give $T_{*}\approx\{7\times 10^{6}\ \rm GeV,7\times 10^{5}\ \rm GeV,7\times 10^{4}\ \rm GeV\}$ and $\beta/H_{*}\approx 1$ for $M_{\rm PBH}=\{10^{3}\ \rm g,10^{6}\ \rm g,10^{9}\ \rm g\}$ respectively. Reducing further the value of $\eta_{\chi}$ would result in too small $\beta/H_{*}$ that is not viable in most scenarios of FOPT. Therefore, the mentioned set of $\eta_{\chi}$ is a good benchmark point as it gives the strongest possible GW signal. The total GW spectrum from the FOPT is shown in Fig. 2 for different PBH’s masses and $f_{\rm PBH}=1$. For an example of individual contributions to the GW spectrum, see Appendix B.
@@
  • The peak frequency is inversely proportional to the mass, which is similar to the inflation case. However, unlike inflation, it also depends on the PBH’s final population (see Eqs. 40 and 23). We observe that, although the magnitude of the GW peak depends on parameter choice of the underlying FOPT, the peak frequency is a robust prediction when we look for a specific mass and population of PBH.
@@
In the left panel of Fig. 3, we show $\Omega_{\rm GW,peak}$ as a function of $\eta_{\chi}$ for different PBH masses. Other fixed benchmark parameters are the same as before and are shown on the top of the panel. The light portion of the curves cannot be detected by current proposed experiments [^8] These just include some major proposals such as LISA, BBO, ET. Of course, if it happens that we can have more experiments with greater sensitivity, then more regions of the parameter space can be probed.. For each value of $\eta_{\chi}$ determined from the left panel, we can infer the corresponding values of $\beta/H_{*}$ and $T_{*}$, as shown in the right panel of Fig. 3. Therefore, a detection of such GW could be an indication of burdened PBHs as DM scenario and could also offer insights into the underlying details of FOPT.
Some comments are in order regarding the prospect for the required value of $\eta_{\chi}$. As mentioned earlier, the asymmetry parameter for the dark sector has no lower bound and therefore it can, in principle, take arbitrarily small value. For example, in a model involving right-handed neutrinos decaying into $\chi$ and some scalar $S$ discussed in Hong:2020est , $\eta_{\chi}\sim 10^{-16}$ can be achieved if $M_{S}\approx 0.9988M_{\nu}$. Such small value of $\eta_{\chi}$ is only necessary for $M_{\rm PBH}=10^{3}\rm g$ in order to have a sufficiently strong GW signal. For other benchmark such as $M_{\rm PBH}=10^{9}\rm g$, $\eta_{\chi}$ can be as large as $10^{-12}$ and the signal is still strong enough to be detected (see Fig. 3). Our generic phenomenological approach is therefore a useful guideline for model builders.
For burdened PBHs with mass inside the DM window, they are not evaporating now and their partial evaporation has already been completed before BBN. Therefore, GW is the only way to probe these PBHs. However, there are two situations where we can have additional complementary probes of the burdened PBHs-as-DM scenario:
(1) If PBHs are locked in binary systems, they can merge and form new PBHs. These fresh PBHs are no longer subject to memory burden and therefore can emit some high-energy gamma rays or neutrinos that should be detected by current experiments Zantedeschi:2024ram . The evaporation will be suppressed again after the half-decay time and the PBHs become stabilized.
(2) If PBHs collide with neutron stars (NS), they can produce the fast radio bursts (FRBs) Kainulainen:2021rbg . FRBs are short pulses in the radio frequency spectrum whose origin is currently unknown. The first detection of such FRBs was reported in Lorimer:2007qn based on the data from the Parkes radio telescope. A possible explanation for this signal is that, when a PBH collide with a NS, a FRB could be produced due to the release of the magnetic field energy at the final stage of the swallowing process. The authors of Kainulainen:2021rbg found that PBHs with mass below $10^{22}\rm g$ could reproduce the observed FRB signal, within an order of magnitude estimate. Besides the usual asteroid-mass window, this result also allows for the possibility of burdened PBHs as DM. Now, because the peak frequency of GW is inversely proportional to the mass ($f_{\rm peak,sw}\propto M_{\rm PBH}^{-1/3}$ from Eqs. 40 and 23), the asteroid-mass window would give a peak at much lower frequency outside of the detectors’ sensitivity regions. The most promising case is $M_{\rm PBH}=10^{17}\rm g$ that gives $f_{\rm peak,sw}\sim 10^{-5}\ \rm Hz$, which is barely at the edge of LISA. Therefore, if the observed FRB indeed came from PBH-NS collision and the GW is detected in future experiments at the right frequency that we showed here, then we can say with high probability that burdened PBHs made up the DM.
@@ IV Conclusion
Appendix A: Redshift of GW spectrum in generic cosmological backgrounds